OpenStudy (anonymous):

A rectangle has a base 5 cm and a height of (x + 2) cm. Its area is 20cm^2. What is the value of x?

OpenStudy (anonymous):

Area of rectangle= Base 1 * Base 2., so it's area is 5(x+2) = 20. Divide both sides by 5; x+2= 4, then subtract 2 from both sides to get x

OpenStudy (campbell_st):

well Area = base x height so 20 = 5*(x + 2) divide both sides of the equation by 5 then subtract 2 from both sides

OpenStudy (anonymous):

So x= 4-2=2

OpenStudy (anonymous):

Thank you, again cx Thank you as well @campbell_st ♥

OpenStudy (anonymous):

Just glad to help :)